# Diffraction-Aided Wireless Positioning

### v2

@@ I Introduction
  • O2I Signal Propagation Insights: We argue that diffraction from window edges plays a key role in outdoor-to-indoor (O2I) signal propagation from a transmitter located outside a building to an indoor receiver. To investigate this further we develop a simplified model of a building. Applying the Geometrical Theory of Diffraction to this building model, we predict the presence of two diffraction multipath components (MPC’s) resulting from diffraction from the upper and lower edges of the building windows. Further, by controlling the transmit side polarization we can control the relative power in the two diffraction MPCs and this leads to a single dominating diffraction-based MPC. Lastly, we also predict that this dominating diffraction MPC in the absence of a direct Euclidean path typically is the first arriving path.
@@ III-A Simplified Model of a Building
Our simplified model of the building begins by assuming that every floor of the building can be modeled as two half-planes separated by a vertical distance $w$ corresponding to the vertical dimension of the window, as shown in Fig.,1(b). Consider a node that is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. The transmit signal originates from the $j^{\text{th}}$ anchor at a known location $\bm{X}_{a,j}=[x_{a,j},y_{a,j},z_{a,j}]^{T}$ below the building floor where the node is located, such that there is NLoS signal propagation. At the node location, we would obtain two diffraction MPCs resulting from the diffraction from the upper and lower edges of the window located on the same floor of the building. Note that our model does not contain vertical pillars that separate windows on the same floor. This is because, applying the lemma 1 for the vertical edges located on the vertical pillars, we would obtain $\lambda>1$ or $\lambda<0$ when solving the corresponding quadratic equation. This implies that the diffraction point $\bm{Q_{e}}$ lies beyond the endpoints of the vertical edges. In this case, the diffraction would be from the corner where the window’s horizontal and vertical edges meet. This phenomenon is called corner diffraction. In our work, we ignore it because it is known to be significantly weaker compared to edge diffraction [39].
@@ III-B Relative power of the two diffraction MPCs
In Fig. 3, we plot the approximate power ratio expression of the two diffraction MPCs in eq. (LABEL:eq_power_ratio) as a function of the elevation angle $\theta$ to the anchor. This is compared with the exact power ratio which is calculated without Assumption 2 and depends on the node position within the building and shows good agreement. Note, $\theta=90$° represents the building floor at the same height as the UAV-anchor. Now, as we go towards higher floors, $\theta$ decreases and for $\theta<60$°, i.e., for building floors located above the height of the UAV-anchor, the diffraction MPC power ratio is at least 5dB. Therefore, we can conclude that for the locations of the nodes on the upper floors of the building, the diffraction MPC of the upper edge dominates the diffraction MPC of the lower edge, leading to a single dominating diffraction MPC.
@@ IV-B Fisher Information analysis of the 3D position estimate
In this section, we examine the conditions under which we can perform 3D NLoS localization using the ToF-based ranging estimates corresponding to the dominating diffraction MPC using Fisher Information analysis and then calculate the lower bound to the 3D position error. In Fig. 4, we have ‘$M$’ UAV anchors in front of the building. The node is located on an arbitary building floor which has two diffracting edges separated by a vertical distance $w$. We assume that the origin is located in the middle bottom of the front wall of the building at the intersection of the front wall with the ground, the x axis oriented along the diffracting edges and the z-axis pointed along the vertical direction as in Fig.,4.
@@
Since we assume the noise in our range measurements is independent across the $M$ anchors, $\bm{J}_{\tau_{0}}=diag(\lambda_{0},\cdots,\lambda_{M-1})$ is a diagonal matrix where $\lambda_{j}=\frac{1}{8\pi^{2}\beta^{2}SNR_{j}}$ represents the inverse of the CRLB of the path delay estimate of the first arriving path between the $j^{th}$ anchor and the node. The Jacobian $\bm{J}\in\mathbb{R}^{M\times 3}$ is the transform from the ranging measurement space $\in\mathbb{R}^{M}$ to the unknown parameter to be estimated space $\in\mathbb{R}^{3}$ and is given by
$\bm{J}=\begin{bmatrix}\frac{\partial{p_{0}(\bm{\alpha})}}{\partial x_{n}}&\frac{\partial{p_{0}(\bm{\alpha})}}{\partial y_{n}}&\frac{\partial{p_{0}(\bm{\alpha})}}{\partial z_{n}}\\ \vdots&\vdots&\vdots\\ \frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial x_{n}}&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial y_{n}}&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial z_{n}}\\ \end{bmatrix}.$ (20)
@@
  • It is necessary but not sufficient to have at least four anchors. This is because, for $rank(\bm{J})=3$, it is necessary but not sufficient to have $M\geq 3$.
@@ VI Conclusion
Figure 6: Raytracing setup using Remcom’s Wireless Insight [31]. We create a building with floors, ceilings and exterior made out of ITU 28 GHz Concrete and Drywall partitions in the interiors. For numerical results we consider 17485 candidate receiver locations located across five building floors which are NLoS to all four UAV-anchors situated outside the building.
@@
(a) Elevation AoA Boxplot.png
(b) SNR Boxplot.png
@@
$\begin{split}\frac{\partial p}{\partial x_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial x_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(x_{n}-q_{x})(1-\frac{\partial q_{x}}{\partial x_{n}})}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial y_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial y_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial y_{n}}+y_{n}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial z_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial q_{x}}{\partial x_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial x_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial x_{n}}\pm\frac{b\frac{\partial b}{\partial x_{n}}-2c\frac{\partial a}{\partial x_{n}}-2a\frac{\partial c}{\partial x_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial y_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial y_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial y_{n}}\pm\frac{b\frac{\partial b}{\partial y_{n}}-2c\frac{\partial a}{\partial y_{n}}-2a\frac{\partial c}{\partial y_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial z_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial z_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial z_{n}}\pm\frac{b\frac{\partial b}{\partial z_{n}}-2c\frac{\partial a}{\partial z_{n}}-2a\frac{\partial c}{\partial z_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \end{split}$
$\begin{split}\frac{\partial a}{\partial x_{n}}&=0,\;\frac{\partial a}{\partial y_{n}}=2y_{n}(x_{1}-x_{2})^{2},\frac{\partial a}{\partial z_{n}}=2(x_{1}-x_{2})^{2}(z_{n}+\frac{w}{2}),\\ \frac{\partial b}{\partial x_{n}}&=2(x_{1}-x_{2})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial b}{\partial y_{n}}&=4y_{n}(x_{1}-x_{2})(x_{2}-x_{a}),\\ \frac{\partial b}{\partial z_{n}}&=-4(x_{1}-x_{2})(x_{2}-x_{n})(z_{n}+\frac{w}{2}-z_{a}),\\ \frac{\partial c}{\partial x_{n}}&=2(x_{2}-x_{n})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial c}{\partial y_{n}}&=2y_{n}(x_{2}-x_{a})^{2},\frac{\partial c}{\partial z_{n}}=-2(x_{2}-x_{n})^{2}(z_{n}+\frac{w}{2}-z_{a}).\end{split}$ (28)

### v3

@@ I Introduction
  • O2I Signal Propagation Insights: We argue that diffraction from window edges plays a key role in outdoor-to-indoor (O2I) signal propagation from a transmitter located outside a building to an indoor receiver. To investigate this further we develop a simplified model of a building. Applying the Geometrical Theory of Diffraction (GTD) to this building model, we predict the presence of two diffraction multipath components (MPC’s) resulting from diffraction from the upper and lower edges of the building windows. Further, by controlling the transmit side polarization we can control the relative power in the two diffraction MPCs and this leads to a single dominating diffraction-based MPC. Lastly, we also predict that this dominating diffraction MPC in the absence of a direct Euclidean path typically is the first arriving path.
@@ III-A Simplified Model of a Building
Our simplified model of the building begins by assuming that every floor of the building can be modeled as two half-planes separated by a vertical distance $w$ corresponding to the vertical dimension of the window, as shown in Fig.1(b). Consider a node that is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. The transmit signal originates from the $j^{\text{th}}$ anchor at a known location $\bm{X}_{a,j}=[x_{a,j},y_{a,j},z_{a,j}]^{T}$ below the building floor where the node is located, such that there is NLoS signal propagation. At the node location, we would obtain two diffraction MPCs resulting from the diffraction from the upper and lower edges of the window located on the same floor of the building. Note that our model does not contain vertical pillars that separate windows on the same floor. This is because, applying the lemma 1 for the vertical edges located on the vertical pillars, we would obtain $\lambda>1$ or $\lambda<0$ when solving the corresponding quadratic equation. This implies that the diffraction point $\bm{Q_{e}}$ lies beyond the endpoints of the vertical edges. In this case, the diffraction would be from the corner where the window’s horizontal and vertical edges meet. This phenomenon is called corner diffraction. In our work, we ignore it because it is known to be significantly weaker compared to edge diffraction [39].
@@ III-B Relative power of the two diffraction MPCs
In Fig. 3, we plot the approximate power ratio expression of the two diffraction MPCs in eq. (LABEL:eq_power_ratio) as a function of the elevation angle $\theta$ to the anchor. This is compared with the exact power ratio which is calculated without Assumption 2 and depends on the node position within the building and shows good agreement. Note, $\theta=90$° represents the building floor at the same height as the UAV-anchor. Now as we go towards higher floors, $\theta$ decreases and for $\theta<60$°, i.e., for building floors located above the height of the UAV-anchor, the diffraction MPC power ratio is at least 5dB. Therefore, we can conclude that for the locations of the nodes on the upper floors of the building, the diffraction MPC of the upper edge dominates the diffraction MPC of the lower edge, leading to a single dominating diffraction MPC.
@@ IV-B Fisher Information analysis of the 3D position estimate
In this section, we examine the conditions under which we can perform 3D NLoS localization using the ToF-based ranging estimates corresponding to the dominating diffraction MPC using Fisher Information analysis and then calculate the lower bound to the 3D position error. In Fig.4, we have ‘$M$’ UAV anchors in front of the building. The node is located on an arbitary building floor which has two diffracting edges separated by a vertical distance $w$. We assume that the origin is located in the middle bottom of the front wall of the building at the intersection of the front wall with the ground, the x axis oriented along the diffracting edges and the z-axis pointed along the vertical direction as in Fig.4.
@@
Since we assume the noise in our range measurements is independent across the $M$ anchors, $\bm{J}_{\tau_{0}}=diag(\lambda_{0},\cdots,\lambda_{M-1})$ is a diagonal matrix where $\lambda_{j}=8\pi^{2}\beta^{2}SNR_{j}$ represents the inverse of the CRLB of the path delay estimate of the first arriving path between the $j^{th}$ anchor and the node. The Jacobian $\bm{J}\in\mathbb{R}^{3\times M}$ is the transform from the ranging measurement space $\in\mathbb{R}^{M}$ to the unknown parameter to be estimated space $\in\mathbb{R}^{3}$ and is given by
$\bm{J}=\begin{bmatrix}\frac{\partial{p_{0}(\bm{\alpha})}}{\partial x_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial x_{n}}\\ \frac{\partial{p_{0}(\bm{\alpha})}}{\partial y_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial y_{n}}\\ \frac{\partial{p_{0}(\bm{\alpha})}}{\partial z_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial z_{n}}\end{bmatrix}.$ (20)
@@
  • It is necessary but not sufficient to have at least three anchors. This is because, for $rank(\bm{J})=3$, it is necessary but not sufficient to have $M\geq 3$.
@@ VI Conclusion
Figure 6: Raytracing setup using Remcom’s Wireless Insight [31]. We create a building with floors, ceilings and exterior made out of ITU 28 GHz Concrete and Drywall partitions in the interiors. For numerical results we consider 17,485 candidate receiver locations located across five building floors which are NLoS to all four UAV-anchors situated outside the building.
@@
(a) Elevation AoA Boxplot.
(b) SNR Boxplot.
@@
$\begin{split}\frac{\partial p}{\partial x_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial x_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(x_{n}-q_{x})(1-\frac{\partial q_{x}}{\partial x_{n}})}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial y_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial y_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial y_{n}}+y_{n}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial p}{\partial z_{n}}&=\frac{(q_{x}-x_{a})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{a}-z_{n}-0.5w)^{2}}}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{(q_{x}-x_{n})\frac{\partial q_{x}}{\partial z_{n}}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}},\\ \frac{\partial q_{x}}{\partial x_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial x_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial x_{n}}\pm\frac{b\frac{\partial b}{\partial x_{n}}-2c\frac{\partial a}{\partial x_{n}}-2a\frac{\partial c}{\partial x_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial y_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial y_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial y_{n}}\pm\frac{b\frac{\partial b}{\partial y_{n}}-2c\frac{\partial a}{\partial y_{n}}-2a\frac{\partial c}{\partial y_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial q_{x}}{\partial z_{n}}&=\frac{(x_{1}-x_{2})}{2a}\left[\frac{\partial a}{\partial z_{n}}\left(\frac{(b\mp\sqrt{b^{2}-4ac}}{a}\right)\right.\\ &\left.\;\;\;\;\;\;\;\;\;\;\;\;+\left(-\frac{\partial b}{\partial z_{n}}\pm\frac{b\frac{\partial b}{\partial z_{n}}-2c\frac{\partial a}{\partial z_{n}}-2a\frac{\partial c}{\partial z_{n}}}{\sqrt{b^{2}-4ac}}\right)\right],\\ \frac{\partial a}{\partial x_{n}}&=0,\;\frac{\partial a}{\partial y_{n}}=2y_{n}(x_{1}-x_{2})^{2},\frac{\partial a}{\partial z_{n}}=2(x_{1}-x_{2})^{2}(z_{n}+\frac{w}{2}),\\ \frac{\partial b}{\partial x_{n}}&=2(x_{1}-x_{2})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \end{split}$
$\begin{split}\frac{\partial b}{\partial y_{n}}&=4y_{n}(x_{1}-x_{2})(x_{2}-x_{a}),\\ \frac{\partial b}{\partial z_{n}}&=-4(x_{1}-x_{2})(x_{2}-x_{n})(z_{n}+\frac{w}{2}-z_{a}),\\ \frac{\partial c}{\partial x_{n}}&=2(x_{2}-x_{n})\left((z_{1}-z_{a})^{2}+y_{a}^{2}\right),\\ \frac{\partial c}{\partial y_{n}}&=2y_{n}(x_{2}-x_{a})^{2},\frac{\partial c}{\partial z_{n}}=-2(x_{2}-x_{n})^{2}(z_{n}+\frac{w}{2}-z_{a}).\end{split}$ (28)